# Large Language Models for Limited Noisy Data: A Gravitational Wave Identification Study

### v1

@@ Abstract
This work investigates whether large language models (LLMs) offer advantages over traditional neural networks for astronomical data processing, in regimes with non-Gaussian, non-stationary noise and limited labeled samples. Gravitational wave observations provide an suitable test case, using only 90 LIGO events, finetuned LLMs achieve 97.4% accuracy for identifying signals. Further experiments show that, in contrast to traditional networks that rely on large simulated datasets, additional simulated samples do not improve LLM performance, while scaling studies reveal predictable gains with increasing model size and dataset size. These results indicate that LLMs can extract discriminative structure directly from observational data and provide an efficient assessment for gravitational wave identification. The same strategy may extend to other astronomical domains with similar noise properties, such as radio or pulsar observations.
1 Introduction
@@
This article is structured as follows. Section 2 outlines the motivation for this work. Building on how LLMs process data and on our previous finetuning experience, we hypothesize that LLMs may handle non-Gaussian and non-stationary noise more effectively than traditional neural networks and also adapt well when only limited training data are available. This motivates a direct experimental evaluation. Section 3 presents our observational data experiment, including dataset construction, model configuration, training, and classification results. Section 4 extends the analysis by examining the influence of simulated datasets, model size, and dataset size. Section 5 discusses how similar data characteristics arise in other astronomical domains. Finally, Section 6 summarizes the main findings of this study.
2 Motivation & Prediction
2.1 Non-Gaussian and Non-Stationary Noise
@@
producing a smooth, monotonic chirp track in the time–frequency plane. Glitches, lacking long-range coherence, occupy short temporal or spectral intervals.
@@
2.2 Token-Based LLMs Match the Noise Structure
@@
where $I(x,y)$ is the input image, $w_{k}$ is the local kernel, $b_{k}$ is the bias, $(u,v)$ is the output location, and $\sigma(\cdot)$ is a nonlinear activation. Since each activation is dominated by a small neighborhood of pixels, any strong transient spike inside that neighborhood can overwhelm the representation, causing CNNs to misinterpret glitches as salient features. Meanwhile, the global waveform morphology, distributed across many time–frequency bins, may receive little weight.
@@
where $E(\cdot)$ is a learned embedding function. Each token $x_{i}$ therefore summarizes the overall pattern inside patch $P_{i}$, rather than the value of any single pixel. Localized noise bursts typically affect only a few pixels inside a patch and therefore only weakly perturb its embedding, while a coherent waveform leaves correlated structure across multiple patches.
@@
3 Testing LLMs by Gravitational Wave
3.1 Gravitational Wave Dataset Construction
@@ 3.2 Model Architecture
3.3 Model Training
Finetuning of the Meta-Llama-3-8B-Instruct model is conducted on three RTX 3090 GPUs. To accommodate memory constraints, the batch size was set to 1 per GPU, with gradient accumulation of 8 steps to simulate a larger effective batch size. A relatively high learning rate of $5.0\times 10^{-4}$ was used to accelerate convergence, combined with a weight decay of $1.0\times 10^{-3}$ to prevent overfitting. The training process employed a cosine annealing scheduler (Loshchilov & Hutter, 2016), which provides smooth and gradual learning rate decay and demonstrated stable performance across experiments.
@@ 3.3 Model Training
Model training uses the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of $10^{-5}$. The objective function is the binary cross-entropy loss,
@@
3.4 Results on Observational Data
@@
On the held-out test set, the finetuned model achieves balanced and stable classification performance. As illustrated in Figure 2, the recall for both the positive class (segments containing gravitational wave signals) and the negative class (noise only segments) reaches 97.4%, corresponding to a misclassification rate of approximately 2.6% per class. These results demonstrate that, in the presence of strongly non-Gaussian and non-stationary detector noise, the model maintains high detection efficiency while keeping the false-alarm rate low. The nearly symmetric performance between the two classes further indicates that the finetuning procedure does not bias the model toward any particular outcome.
@@
Figure 2: Identification performance of the finetuned LLM on LIGO observational data. The model, trained on only 90 events, achieves a recall of 97.4% for both signal segments and noise segments. The misclassification rate of 2.6% per class demonstrates that the model maintains balanced and reliable performance despite the strongly non-Gaussian and non-stationary noise present in detector data.
4 Further Investigation
4.1 Does Large Simulated Dataset Improve LLM Performance?
@@
4.2 How Does Model Size Improve LLMs Performance?
@@
4.3 How does Dataset Size Improve LLMs Performance ?
@@
5 Applicability to Other Astronomical Domains
@@
These conditions are not unique to gravitational wave detectors. Several astronomical domains produce time–frequency or time–channel data in which coherent astrophysical structures must be extracted from complex, transient dominated noise. Examples include radio pulsar and fast radio burst searches, dynamic spectra from low frequency interferometers, X-ray timing of accreting neutron stars and black holes, and high energy transient monitors (Vaughan, 2013; Linares et al., 2013; Salvo & Sanna, 2021; Hurley-Walker et al., 2022; Dong & Hallinan, 2022; Li et al., 2023; Anumarlapudi et al., 2025). In many of these settings, the signal of interest forms a smooth or slowly evolving pattern, while the noise sometimes contains short lived, non-stationary, and instrument specific features.

### v2

@@ Abstract
This work investigates whether large language models (LLMs) offer advantages over traditional neural networks for astronomical data processing, in regimes with non-Gaussian, non-stationary noise and limited labeled samples. Gravitational wave observations provide a suitable test case, using only 90 LIGO events, finetuned LLMs achieve 97.4% accuracy for identifying signals. Further experiments show that, in contrast to traditional networks that rely on large simulated datasets, additional simulated samples do not improve LLM performance, while scaling studies reveal predictable gains with increasing model size and dataset size. These results indicate that LLMs can extract discriminative structure directly from observational data and provide an efficient assessment for gravitational wave identification. The same strategy may extend to other astronomical domains with similar noise properties, such as radio or pulsar observations.
I Introduction
@@
This article is structured as follows. Section II outlines the motivation for this work. Building on how LLMs process data and on our previous finetuning experience, we hypothesize that LLMs may handle non-Gaussian and non-stationary noise more effectively than traditional neural networks and also adapt well when only limited training data are available. This motivates a direct experimental evaluation. Section III presents our observational data experiment, including dataset construction, model configuration, training, and classification results. Section IV extends the analysis by examining the influence of simulated datasets, model size, and dataset size. Section V discusses how similar data characteristics arise in other astronomical domains. Finally, Section VI summarizes the main findings of this study.
II Motivation & Prediction
II.1 Non-Gaussian and Non-Stationary Noise
@@
producing a smooth, monotonic chirp track in the time–frequency plane. Glitches, lacking long-range coherence, occupy short temporal or spectral intervals. Even in the seconds long cases, they often appear as broadband noise blocks, irregularly drifting streaks, or piecewise concatenated structures. They can be only long in duration, but they are not coherent.
@@
II.2 Token-Based LLMs Match the Noise Structure
@@
where $I(x,y)$ is the input image, $w_{k}$ is the local kernel, $b_{k}$ is the bias, $(u,v)$ is the output location, and $\sigma(\cdot)$ is a nonlinear activation. Since each activation is dominated by a small neighborhood of pixels, any strong transient spike inside that neighborhood can overwhelm the representation, causing CNNs to misinterpret glitches as salient features. Meanwhile, the global waveform morphology, distributed across many time–frequency bins, may receive little weight. This effect is stronger for seconds long glitches. A CNN expands its receptive field by stacking layers, and at finite depth it still represents structures on multi second timescales by piecing together local features. If a glitch shows high energy or sharp edges in many local windows, these local responses can accumulate, and the model can easily interpret the result as a signal like pattern. This failure is not because the glitch is too short, but because it is long enough in time to keep activating local convolutional kernels.
@@
where $E(\cdot)$ is a learned embedding function. Each token $x_{i}$ therefore summarizes the overall pattern inside patch $P_{i}$, rather than the value of any single pixel. Localized noise bursts typically affect only a few pixels inside a patch and therefore only weakly perturb its embedding, while a coherent waveform leaves correlated structure across multiple patches. When a transient spans a longer interval, it can influence many patches, but the embeddings still reflect whether neighboring patches participate in a consistent track or vary irregularly across time and frequency.
@@
III Testing LLMs by Gravitational Wave
III.1 Gravitational Wave Dataset Construction
@@ III.2 Model Architecture
III.3 Model Training
Finetuning of the Meta-Llama-3-8B-Instruct model is conducted on three RTX 3090 GPUs. To accommodate memory constraints, the batch size was set to 1 per GPU, with gradient accumulation of 8 steps to simulate a larger effective batch size. AdamW optimizer (Loshchilov & Hutter, 2017) with a relatively high learning rate of $5.0\times 10^{-4}$ was used to accelerate convergence, combined with a weight decay of $1.0\times 10^{-3}$ to prevent overfitting. The training process employed a cosine annealing scheduler (Loshchilov & Hutter, 2016), which provides smooth and gradual learning rate decay and demonstrated stable performance across experiments.
@@ III.3 Model Training
The objective function is the binary cross-entropy loss,
@@
III.4 Results on Observational Data
@@
On the held-out test set, the finetuned model achieves balanced and stable classification performance. As illustrated in Figure 2, the recall for both the positive class (segments containing gravitational wave signals) and the negative class (noise only segments) reaches 97.4%, corresponding to a misclassification rate of approximately 2.6% per class and the entire accuracy of 97.4%. These results demonstrate that, in the presence of strongly non-Gaussian and non-stationary detector noise, the model maintains high detection efficiency while keeping the false-alarm rate low. The nearly symmetric performance between the two classes further indicates that the finetuning procedure does not bias the model toward any particular outcome.
@@
Figure 2: Identification performance of the finetuned LLM on LIGO observational data. The model, trained on only 90 events, achieves a recall of 97.4% for both signal segments and noise segments, and an accuracy of 97.4%. The misclassification rate of 2.6% per class demonstrates that the model maintains balanced and reliable performance despite the strongly non-Gaussian and non-stationary noise present in detector data.
IV Further Investigation
IV.1 Does Large Simulated Dataset Improve LLM Performance?
@@
IV.2 How Does Model Size Improve LLMs Performance?
@@
IV.3 How does Dataset Size Improve LLMs Performance ?
@@
V Applicability to Other Astronomical Domains
@@
These conditions are not unique to gravitational wave detectors. Several astronomical domains produce time-frequency or time–channel data in which coherent astrophysical structures must be extracted from complex, transient dominated noise. Examples include radio pulsar and fast radio burst searches, dynamic spectra from low frequency interferometers, X-ray timing of accreting neutron stars and black holes, and high energy transient monitors (Vaughan, 2013; Linares et al., 2013; Salvo & Sanna, 2021; Hurley-Walker et al., 2022; Dong & Hallinan, 2022; Li et al., 2023; Anumarlapudi et al., 2025). In many of these settings, the signal of interest forms a smooth or slowly evolving pattern, while the noise sometimes contains short lived, non-stationary, and instrument specific features.